# Multi-Prompting Decoder Helps Better Language Understanding

### v1

@@ V-H Effects of Prompts (RQ4)
[…]
TABLE XI: Accuracy using various types of prompts on AG’s news dataset under the 16-shot setting.
@@
        V+I         87.49
        V+E         88.11
        I+E         87.66
        V+I+E         87.92
V-I Effectiveness on Various Types of Prompts (RQ4)
In this section, we further show the effectiveness on various types of prompts, including instruction and example. Specifically, we use the vanilla prompt 1 in Table VIII (denoted as V), instruction (denoted as I), and example (denoted as E, we use the first training data from the AG’s news dataset to construct example prompt) on the AG’s news dataset under the 16-shot setting. Three prompts as shown in Table IX.
Firstly, our method usually achieves performance improvement compared to using a single prompt in Table XI. Specifically, using two prompts exceeds using either prompt individually. For example, V+I exceeds V and I, respectively. It is worth noting that V+E exceeds all results in Figure 4. This shows that using two different types of prompts sometimes yields better results. Secondly, the performance fluctuation of using two prompts is smaller than that of a single prompt. Thirdly, V+I+E does not exceed V+I due to the poor performance of I. This is because RoBERTa-large has not been fine-tuned by instruction-tuning.
V-J Extension to Other PLMs (RQ5)
[…]
TABLE XII: The results of label words expansion.
@@ V-J Extension to Other PLMs (RQ5)
Figure 5: Effects of $\beta$ over 5 runs on two datasets.
V-K Effects of Hyper-Parameter $\beta$
In this section, we explore the effects of $\beta$ on two datasets in Figure 5. Hyper-parameter $\beta$ controls the ratio of OT scores and class scores.
Firstly, our model needs a larger $\beta$ under the 1-shot setting. This is because the amount of data is small and the model relies heavily on class scores for prediction rather than OT scores. Thus, a larger $\beta$ usually achieves better performance under the 1-shot setting. Secondly, as the shot number increases, the effect of $\beta$ on the results diminishes. This is because as the shot number increases, the performance of OT scores is enhanced and relies more on the OT scores for prediction. Thirdly, as the shot number increases, the standard deviation becomes smaller. This shows that the OT scores are more stable when the training size increases.
Figure 6: Effects of $Q$ over 5 runs on two datasets.
V-L Effects of the Number of Prototypes
In this section, we explore the effects of the number of prototypes on two datasets in Figure 6.
Firstly, we can see that using multiple prototypes tends to outperform using one. This shows that using multiple prototypes is effective in our multiple prompts setting. Secondly, the improvement is more significant under the 4-shot and 16-shot. This shows that as the shot number increases, multiple prototypes are more effective. Finally, as the shot number increases, a large $Q$ usually performs better. For example, the model achieves the best performance on two datasets under the 4-shot setting with $Q=4$ and 3. The model achieves the best performance on two datasets under the 4-shot setting with $Q=5$ and 4. This is because when the shot number increases, the model needs a large $Q$ to capture subtle differences between multiple texts of the same class. We simply set $Q=3$ on all datasets and settings to avoid hyper-parameter selection.
TABLE XIII: Performance of two variants to replace optimal transport plan. We run each experiment over 5 random seeds.
Model Setting Average Accuracy
1 4 16
Ours 71.5 75.2 77.8
Uniform Distribution 70.7 74.7 77.1
Cosine Similarity 68.3 73.8 77.4
V-M Visualization of Label Words Expansion
We show the expanded label words in Table XII. We can see that our proposed simple expansion method expands some similar words and some semantically related words. For example, on the three sentiment datasets (i.e., SST2, IMDB, and Yelp), label word ‘bad’ expands some semantically relevant words (i.e., ‘poor’, ‘wrong’, and ‘worst’), and label word ‘great’ expands some relevant words (i.e., ‘good’, ‘perfect’, and ‘powerful’). On the AG’s News dataset, label word ‘politics’ expands ‘democracy’ and ‘history’, and label word ‘technology’ expands ‘software’, ‘science’, and so on. This shows that our proposed label words expansion is effective. While ‘bad’ expands ‘good’, we assign a weight to ‘good’ based on its cosine similarity to the target class word to avoid the effect of noise.
V-N Effectiveness of OT
We use two variants to illustrate the effectiveness of fine-grained matching using OT, including uniform distribution and cosine similarity.
We can see that OT achieves the best performance in Table XIII. This shows that OT can effectively match the relationship between text representations and prototypes. The uniform distribution does not take into account the similarity between representations and prototypes. Cosine similarity only calculates the similarity between two elements and cannot take into account the similarity with other elements.
VI Conclusion
In this paper, we propose the multi-prompting decoder framework for output-side MaaS adaptation on few-shot tasks. To achieve better decoding performance, we design two decoding strategies to decode PLMs’ output hidden states and class scores, respectively. Extensive experiments demonstrate that our method achieves new state-of-the-art results under the few-shot setting on multiple natural language understanding tasks, including sentiment analysis, topic classification, and natural language inference. Besides, our experiments also show that our model is lightweight, efficient, and applicable to PLMs of different architectures (e.g., encoder only, decoder only, and encoder-decoder).

### v2

@@ Appendix F Model Scaling
Appendix G Trade-offs between Training Time and Performance
In this section, we further explore the trade-offs between training time and performance in our method.
Firstly, training time is proportional to the number of adopted prompts from Table 9. This is because we need to feed more training data into PLM. Secondly, when the number of prompts exceeds 3, the improvement becomes marginal. As a result, to balance training time and performance, we use three prompts.
Adopted Prompts Acc Training Time (s)
1 87.44 2.3
1+2 87.82 4.5
1+2+3 87.94 6.7
1+2+3+5 87.98 8.8
1+2+3+4+5 87.97 10.9
Table 9: Accuracy and training time using different prompts on AG’s news dataset under 16-shot setting.
@@ Appendix G Trade-offs between Training Time and Performance
        Vani+Instr         87.49
        Vani+E         88.11
        Instr+E         87.66
        Vani+Instr+E         87.92
Table 10: Accuracy using various types of prompts on AG’s news dataset under the 16-shot setting.
@@
Table 11: The results of label words expansion.
Dataset Template Label Words
SST2 $x$ A [MASK] movie. bad, great
$x$ A [MASK] film.
$x$ A [MASK] piece of work.
Yelp $x$ In summary, it was [MASK]. bad, great
$x$ All in all, it was [MASK].
$x$ A [MASK] review.
IMDB $x$ In summary, it was [MASK]. bad, great
$x$ All in all, it was [MASK].
$x$ In summary, the film was [MASK].
AG’s News [ Topic : [MASK] ] $x_{1}$ $x_{2}$ politics, sports, business, technology
[ Category : [MASK] ] $x_{1}$ $x_{2}$
$x_{1}$ $x_{2}$ The topic is about [MASK].
Yahoo [ Topic : [MASK] ] $x_{1}$ $x_{2}$ society, science, health, education, computers, sports, business, entertainment, family, politics
A [MASK] question : $x_{1}$ $x_{2}$
$x_{1}$ $x_{2}$ The topic is about [MASK].
DBPedia [ Topic : [MASK] ] $x_{1}$ $x_{2}$ company, school, artist, athlete, politics, transportation, building, river, village, animal, plant, album, film, book
$x_{1}$ $x_{2}$ $x_{1}$ is a [MASK].
$x_{1}$ $x_{2}$ In this sentence, $x_{1}$ is a [MASK].
RTE $x_{1}$? [MASK], $x_{2}$ No, Yes
SNLI $x_{1}$. [MASK], $x_{2}$ No, Maybe, Yes
MNLI-m/mm $x_{1}$! [MASK], $x_{2}$ No, Maybe, Yes
[…]
Template Id Template
1 [ Topic : [MASK] ] $x_{1}$ $x_{2}$
2 [ Category : [MASK] ] $x_{1}$ $x_{2}$
3 $x_{1}$ $x_{2}$ The topic is about [MASK].
4 [MASK] Alert Blog Dialogue Diary Accountability $x_{1}$ $x_{2}$
5 A [MASK] news : $x_{1}$ $x_{2}$
Table 13: Five templates used in our experiments. The fourth prompt comes from Deng et al. (2022).
Template Type Template
Vani [ Topic : [MASK] ] $x_{1}$ $x_{2}$
Instr Classify the news articles into the categories of Politics, Sports, Business, and Technology. $x_{1}$ $x_{2}$ This topic is about [MASK].
E Article: Wall St. Bears Claw Back Into the Black. (Reuters), Reuters - Short-sellers, Wall Street’s dwindling of ultra-cynics, are seeing green again. Answer: Business. Article: $x_{1}$ $x_{2}$ Answer: [MASK].
Table 14: Three different types of templates (i.e., vanilla prompt, instruction prompt, and example prompt) used in our experiments.
Appendix H Effectiveness on Various Types of Prompts
In this section, we further show the effectiveness on various types of prompts, including instruction and example. Specifically, we use the vanilla prompt 1 in Figure 4 (denoted as Vani), instruction (denoted as Instr), and example (denoted as E, we use the first training data from the AG‘s news dataset to construct example prompt) on the AG’s news dataset under the 16-shot setting. Three prompts as shown in Table 14.
Firstly, our method usually achieves performance improvement compared to using a single prompt in Table 10. Specifically, using two prompts exceeds using either prompt individually. For example, Vani+Instr exceeds Vani and Instr, respectively. It is worth noting that Vani+E exceeds all results in Figure 4. This shows that using two different types of prompts can sometimes yield better results. Secondly, the performance fluctuation of using two prompts is smaller than that of a single prompt. Thirdly, Vani+Instr+E does not exceed Vani+Instr due to the poor performance of Instr. This is because RoBERTa-large has not been fine-tuned by instruction-tuning.
Algorithm 1 The training flow of our method
1:Training set and PLM $\mathcal{M}$.
2:The parameters of linear layer $\mathbf{W}$ and prototypes $\mathbf{R}$
[…]
4:Randomly initialize class prototypes $\bm{R}$
5:Obtain a feature set $\bm{H}$ via PLM $\mathcal{M}$
[…]
7:for epoch $=1,2,\dots,T$ do
8:     Sample a data $x_{i}$ and use linear layer to get the text representations $\bm{V}_{i}$
[…]
10:     for $k=1,2,\dots,N$ do
11:         Calculate the cost matrix $\bm{C}_{k}=\bm{1}-\bm{V}\bm{R}_{k}^{\top}$ of each class
12:         for $t_{in}=1,2,\dots,T_{in}$ do
13:              $Z=\exp(-\bm{C}/\lambda)$
14:              $\bm{u}^{(t_{in})}=\bm{u}/(Z\bm{v}^{(t_{in}-1)})$
15:              $\bm{v}^{(t_{in})}=\bm{v}/(Z^{\top}\bm{u}^{(t_{in})})$
16:              $\Delta_{v}=\sum|\bm{v}^{(t_{in})}-\bm{v}^{(t_{in}-1)}|/N$
17:              if $\Delta_{v}<\delta$ then
18:                  break
19:              end if
20:         end for
21:         Obtain optimal transport plan as $\bm{T}^{*}_{k}=\text{diag}(\bm{u}^{(t)})\exp(-\bm{C}_{k}/\lambda)\text{diag}(\bm{v}^{(t)})$
22:         Calculate the OT scores by Eq. (2)
23:     end for
24:     Update the parameters with cross-entropy loss by Eq. (3)
25:end for